What is Ohm's law

Grade:12th pass

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear amarnath

To make a current flow through a resistance there must be a voltage across that resistance. Ohm's Law shows the relationship between the voltage (V), current (I) and resistance (R). It can be written in three ways:

V=IR    or  I=V/R    or    R=V/I

where: V = voltage in volts (V)
I  = current in amps (A)
R = resistance in ohms ()
or: V = voltage in volts (V)
I  = current in milliamps (mA)
R = resistance in kilohms (k)

For most electronic circuits the amp is too large and the ohm is too small, so we often measure current in milliamps (mA) and resistance in kilohms (k). 1 mA = 0.001 A and 1 k = 1000 .

The Ohm's Law equations work if you use V, A and , or if you use V, mA and k. You must not mix these sets of units in the equations so you may need to convert between mA and A or k and .



Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin